Lean Accounting

Westgate Inc. uses a lean manufacturing strategy to manufacture DVR (digital video recorder) players. The company manufactures DVR players through a single product cell. The budgeted conversion cost for the year is $924,000 for 2,200 production hours. Each unit requires 10 minutes of cell process time. During March, 880 DVR players were manufactured in the cell. The materials cost per unit is $77. The following summary transactions took place during March:

  1. Materials were purchased for March production.
  2. Conversion costs were applied to production.
  3. 880 DVR players were assembled and placed in finished goods.
  4. 840 DVR players were sold for $260 per unit.

a. Determine the budgeted cell conversion cost per hour. If required, round to the nearest dollar.
$ per hour

b. Determine the budgeted cell conversion cost per unit. If required, round to the nearest dollar.
$ per unit

c. Journalize the summary transactions (1)–(4) for March.

1.
2.
3.
4. Sale
4. Cost
Answer #1
a
Budgeted cell conversion cost per hour 420 per hour =924000/2200
b
Budgeted cell conversion cost per unit 70.00 per unit =420*(10/60)
c
Raw and in process inventory 67760 =880*77
     Accounts Payable 67760
Raw and in process inventory 61600 =880*70
       Conversion costs 61600
Finished goods inventory 129360 =880*(77+70)
       Raw and in process inventory 129360
Accounts Receivable 218400 =840*260
       Sales 218400
Cost of goods sold 123480 =840*(77+70)
        Finished goods inventory 123480